OpenStudy (anonymous):

I need the square root of x+2=X...

OpenStudy (tanya123):

sqrt(x+2) = X

OpenStudy (anonymous):

\[\sqrt{x+2}=x\] just like that

OpenStudy (anonymous):

yes what? I need to solve that.

OpenStudy (tanya123):

x = X^2-2

OpenStudy (tanya123):

So, x = -2

OpenStudy (anonymous):

wouldn't it be \[x+2=x ^{2}\] ?

OpenStudy (anonymous):

Oh, I think I see what your saying

OpenStudy (anonymous):

just a sec

OpenStudy (tanya123):

Should look like this, http://prntscr.com/44rhng I'm not able to make it in the laTex form

OpenStudy (anonymous):

but i'd have to preform the operation \[x ^{2}\]/x while subtracting 2

OpenStudy (tanya123):

Hold on, he'll explain you @TwinYang

OpenStudy (anonymous):

as in, Simultaneously?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

steps.

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

Do I need to type new quiestions in as that?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

Wait, why did you multiply all terms in the equation by -1

OpenStudy (anonymous):

because we used the polynomial method to get it

OpenStudy (anonymous):

polynomial method. Got it. I'll need to write a note on that. I didn't pay much attention this week, and i'm suffering for it. Thank you.

OpenStudy (anonymous):

welcome